OpenStudy (anonymous):

Find all values of t in the interval [0, 2 π] that satisfy cos^2 t-4cost+3=0 . (Hint: Substitute x = cos t, factor the result, find the values of x that solve the equation, then put cos t back in for x.)

OpenStudy (anonymous):

Comment: The equation should be cos2 t - 4cos t + 3 = 0. Rewrite this as x2 - 4x + 3 = 0 and factor to solve for x. Now going backward cos t = the numbers you got for x. You should be able to find the correct answers from the Unit Circle.

OpenStudy (anonymous):

I solved x2 - 4x + 3 = 0and got x=1 and x=3

OpenStudy (anonymous):

good work, now, \[\cos t=1\qquad\qquad\qquad\cos t=3\]

OpenStudy (anonymous):

for what values of "t" is cos(t)=1??

OpenStudy (anonymous):

I got 2piN1

OpenStudy (anonymous):

I do not understand that

OpenStudy (anonymous):

When it says backwards, is it wanting the cos^-1?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

your domain is a closed interval.. so when t = 0 and when t = \(2\pi\) \(\cos(t)=1\) correct?

OpenStudy (anonymous):

I got 0.5403.

OpenStudy (anonymous):

how is that?

OpenStudy (anonymous):

I did cos(1)

OpenStudy (anonymous):

no sine \(\cos(t)=1\) we do \(t=\cos^{-1}(1)\) get it?

OpenStudy (anonymous):

now read my previous setps

OpenStudy (anonymous):

I got t=0.

OpenStudy (anonymous):

can there be any other possibilities?

OpenStudy (anonymous):

2pi?

OpenStudy (anonymous):

There is no solution for t=cos^-1(3)

OpenStudy (anonymous):

good. all other values will not fit in our domain. so, these are the two solutions for the first one now, the last one t=cos^-1(3) can never happen

OpenStudy (anonymous):

good job.

OpenStudy (anonymous):

Thank you for the help!

OpenStudy (anonymous):

yw